# HiLAB: A Hybrid Inverse-Design Framework

### v1

@@ 3.2 Multispectral Characterization System and Results
The characterization system is schematically shown in Figure 7(d). A SuperK FIANIUM broadband source (NKT Photonics) precisely generates red, green, and blue wavelengths at 660 nm, 550 nm, and 470 nm, respectively, forming a multi-wavelength beam. This beam is first linearly polarized by passing through a linear polarizer.
Figure 7: (a) Refractive-index profile of the optimized metasurface unit cell. (b) SEM of the fabricated metasurface deflector. The inset (red box) highlights the subwavelength corrugations critical for broadband operation. Scale bars: $5\,\mu\mathrm{m}$ (main image) and $1\,\mu\mathrm{m}$ (inset). (c) Experimentally measured deflection efficiency as a function of output angle for three design wavelengths: 470 nm (blue), 550 nm (green), and 660 nm (red). The prominent efficiency peaks near $\theta=41.3^{\circ}$ confirm the metasurface’s broadband and angle-specific beam deflection performance. The measured efficiencies agree well with theoretical simulations (25.0% for red , 24.2% for green, and 24.1% for blue). (d) Schematic of the broadband optical characterization setup. A collimated beam from a broadband source passes through a linear polarizer and imaging optics, illuminates the object and metasurface, and is then focused onto a CCD detector. The system enables angle-resolved measurements, with the detection angle $\theta=41.3\degree$ indicated. (e) Optical characterization using a USAF 1951 resolution target under illumination at the same three wavelengths and their combinations. Individual and combined RGB (red, green, blue) channel images validate achromatic beam deflection across the visible spectrum.
@@
To measure the deflection efficiency at each wavelength, the power of the incident plane-wave at that wavelength is measured by a detector placed right before the metasurface. The output power is then measured as a function of the deflected angle and divided by the total incident power to obtain the deflection efficiency at the given wavelength. Figure 7(c) shows the measured deflection efficiency as a function of the angle for the three wavelengths. It is clearly seen that all beams deflect at $41.3^{\circ}$ with similar efficiencies. The experimentally measured efficiencies—22.1% for red, 20.5% for green, and 17.7% for blue—demonstrate good agreement with theoretical results and confirm robust broadband deflection performance. This represents a major improvement over previously published results [28], both in terms of the uniformity of efficiency over a wide spectral range and the absolute diffraction efficiencies. These results highlight the clear advantage of HiLAB over conventional TO in finding an optimal design. The slightly lower experimental efficiency observed for blue light is primarily attributed to material absorption. While the loss of TiO2 at visible wavelengths is generally low, its extinction coefficient $k$ becomes slightly elevated at shorter wavelengths, resulting in increased absorption for blue light. This wavelength-dependent behavior is confirmed by the measured complex refractive index of a TiO2 film, shown in Supplementary Figure S1, where both $n$ and $k$ are plotted as functions of wavelength. In another experiment, the polarized light illuminates a standard USAF 1951 resolution test target and subsequently passes through an objective lens. The beam then illuminates the fabricated achromatic beam deflector (metasurface), which redirects it toward an imaging arm mounted on a large-area rotating breadboard, enabling precise angular scanning of the output beam. Finally, a tube lens relays the image onto a CCD camera, capturing the test target pattern with spectral and polarization modifications introduced by the metasurface. The output pattern of the metasurface at the desired wavelength is imaged under illumination with the patterned light at different input light wavelengths and their combinations. The results, shown in Figure 7(e), clearly demonstrate the ability of the fabricated device in deflecting different colors at the same angle with almost the same efficiency.
@@ 3.3 Discussions
The key aspects of HiLAB are the use of conventional gradient-based optimization techniques in a low-dimensional space and using BO to widen the local search region and avoid early convergence to local optima. The resulting reduction in computation requirements and the ability to search for a close-to-global optimum (or a strong local optimum) make it useful for the design of large-scale metasurfaces and addition of multi-functional features. Further optimization and refinement of these individual parts can result in dramatic improvement of the performance of HiLAB and enable its application for a wide range of nanophotonic designs that are not currently possible with existing approaches. Nevertheless, despite being at the early stage of development, the theoretical and experimental results presented in this paper clearly demonstrate the unique features of HiLAB and pinpoint the importance of and the need for using hybrid inverse-design techniques that scale up our design capabilities to larger structures while avoiding weak local optima. Further improvements of such approaches are needed to impose more constraints on the optimization process. Example constraints of interest in nanophotonics include: 1) low sensitivity to fabrication imperfections, 2) increasing the minimum required feature sizes, 3) low sensitivity to environmental changes and alignment errors, and so on. Our proposed hybrid inverse-design framework, HiLAB, is highly general and readily adaptable across a wide variety of photonic inverse design problems. Specifically, HiLAB can utilize established gradient-based optimization methods and commercial software tools [44, 7, 45, 46, 47, 32, 48] to efficiently generate targeted training datasets. These datasets enable effective exploration of the design space through subsequent data-driven approaches.
@@ Conclusion
In this paper, we introduced HiLAB, a novel hybrid inverse-design framework that effectively combines partial TO, generative AI via a ViT-based VAE, and BO to efficiently address complex multi-functional photonic device design challenges. By truncating traditional adjoint-driven TO at early stages and systematically varying key physical parameters, we generate diverse and computationally affordable datasets that capture the critical structural features necessary for high-performance device design. The trained VAE encodes these partially optimized structures into a compact latent space, significantly reducing computational demands for subsequent design exploration. By leveraging BO within this learned latent space, HiLAB systematically guides the search towards global optima (or strong local optima) while simultaneously optimizing physical hyperparameters such as device thickness and period. We demonstrated the effectiveness of HiLAB by designing an achromatic beam deflector achieving balanced diffraction efficiencies, theoretically exceeding 24% across red, green, and blue wavelengths—substantially outperforming similar devices designed by prior conventional TO methods and significantly reducing chromatic aberrations. Crucially, the generality and flexibility of HiLAB mean it can be easily adapted for different design targets, constraints, or fabrication tolerances without necessitating extensive additional simulations or retraining. By providing a systematic, scalable, and efficient inverse-design methodology, HiLAB holds considerable promise for accelerating the development of robust, high-performance photonic and metamaterial devices across diverse application domains.

### v2

@@ 3.2 Multispectral Characterization System and Results
The characterization system is schematically shown in Figure 7(d). A SuperK FIANIUM broadband source (NKT Photonics) precisely generates red, green, and blue wavelengths at 660 nm, 550 nm, and 470 nm, respectively, forming a multi-wavelength beam. This beam is first linearly polarized by passing through a linear polarizer. To measure the deflection efficiency at each wavelength, the power of the incident plane-wave at that wavelength is measured by a detector placed right before the metasurface. The output power is then measured as a function of the deflected angle and divided by the total incident power to obtain the deflection efficiency at the given wavelength. Figure 7(c) shows the measured deflection efficiency as a function of the angle for the three wavelengths. It is clearly seen that all beams deflect at $41.3^{\circ}$ with similar efficiencies. The experimentally measured efficiencies—22.1% for red, 20.5% for green, and 17.7% for blue—demonstrate good agreement with theoretical results and confirm robust broadband deflection performance. This represents a major improvement over previously published results [28], both in terms of the uniformity of efficiency over a wide spectral range and the absolute diffraction efficiencies. These results highlight the clear advantage of HiLAB over conventional TO in finding an optimal design. The slightly lower experimental efficiency observed for blue light is primarily attributed to material absorption. While the loss of TiO2 at visible wavelengths is generally low, its extinction coefficient $k$ becomes slightly elevated at shorter wavelengths, resulting in increased absorption for blue light. This wavelength-dependent behavior is confirmed by the measured complex refractive index of a TiO2 film, shown in Supplementary Figure S4, where both $n$ and $k$ are plotted as functions of wavelength. In another experiment, the polarized light illuminates a standard USAF 1951 resolution test target and subsequently passes through an objective lens. The beam then illuminates the fabricated achromatic beam deflector (metasurface), which redirects it toward an imaging arm mounted on a large-area rotating breadboard, enabling precise angular scanning of the output beam. Finally, a tube lens relays the image onto a CCD camera, capturing the test target pattern with spectral and polarization modifications introduced by the metasurface. The output pattern of the metasurface at the desired wavelength is imaged under illumination with the patterned light at different input light wavelengths and their combinations. The results, shown in Figure 7(e), clearly demonstrate the ability of the fabricated device in deflecting different colors at the same angle with almost the same efficiency.
Figure 7: (a) Refractive-index profile of the optimized metasurface unit cell. (b) SEM of the fabricated metasurface deflector. The inset (red box) highlights the subwavelength corrugations critical for broadband operation. Scale bars: $5\,\mu\mathrm{m}$ (main image) and $1\,\mu\mathrm{m}$ (inset). (c) Experimentally measured deflection efficiency as a function of output angle for three design wavelengths: 470 nm (blue), 550 nm (green), and 660 nm (red). The prominent efficiency peaks near $\theta=41.3^{\circ}$ confirm the metasurface’s broadband and angle-specific beam deflection performance. The measured efficiencies agree well with theoretical simulations (25.0% for red , 24.2% for green, and 24.1% for blue). (d) Schematic of the broadband optical characterization setup. A collimated beam from a broadband source passes through a linear polarizer and imaging optics, illuminates the object and metasurface, and is then focused onto a CCD detector. The system enables angle-resolved measurements, with the detection angle $\theta=41.3^{\circ}$ indicated. (e) Optical characterization using a USAF 1951 resolution target under illumination at the same three wavelengths and their combinations. Individual and combined RGB (red, green, blue) channel images validate achromatic beam deflection across the visible spectrum.
@@ 3.3 Discussions
The key aspects of HiLAB are the use of conventional gradient-based optimization techniques in a low-dimensional space and using BO to widen the local search region and avoid early convergence to local optima. The resulting reduction in computation requirements and the ability to search for a close-to-global optimum (or a strong local optimum) make it useful for the design of large-scale metasurfaces and the addition of multifunctional features. Indeed, this focus on complex functionality [50] aligns with a broader trend in the field, as recent works at the intersection of deep learning and TO using hybrid inverse-design frameworks further highlight, demonstrating intelligent [51] and adaptable nanophotonic systems [52]. These studies underscore the growing need for hybrid frameworks that combine the expressivity of AI with the rigor of physics-based optimization. Further optimization and refinement of these individual parts can result in dramatic improvement of the performance of HiLAB and enable its application for a wide range of nanophotonic designs that are not currently possible with existing approaches. Nevertheless, despite being at the early stage of development, the theoretical and experimental results presented in this paper clearly demonstrate the unique features of HiLAB and pinpoint the importance of and the need for using hybrid inverse-design techniques that scale up our design capabilities to larger structures while avoiding weak local optima. Further improvements of such approaches are needed to impose more constraints on the optimization process. Example constraints of interest in nanophotonics include: 1) low sensitivity to fabrication imperfections, 2) increasing the minimum required feature sizes, 3) low sensitivity to environmental changes and alignment errors, and so on. Our proposed hybrid inverse-design framework, HiLAB, is highly general and readily adaptable across a wide variety of photonic inverse design problems. Specifically, HiLAB can utilize established gradient-based optimization methods and commercial software tools [53, 7, 54, 55, 56, 32, 57] to efficiently generate targeted training datasets. These datasets enable effective exploration of the design space through subsequent data-driven approaches.
The transferability of the trained VAE [58] is a powerful feature of the HiLAB framework, aligning with the development of general design platforms that can adapt to new objectives without complete retraining [59]. In our approach, the VAE learns a physics-agnostic representation of the device geometry [60], leaving the task of solving the physics to the numerical solver. This learned representation can therefore be reused without retraining for other applications involving the same device architecture. The usefulness of this transferred representation will depend on how well the learned geometrical insights apply to the new problem. We have demonstrated this by transferring the latent representation from a VAE trained for a 41.3 beam deflection to a new target of 30 deflection angle without retraining. The transferred representation yielded a comparable FoM of 0.20, confirming that the learned geometrical features are indeed transferable to closely related tasks. It remains an open question whether this representation will remain equally effective when the design objective changes more drastically (e.g., from a beam deflector to a broadband absorber). In such cases, even if direct transfer is less effective, the pretrained representation may still be valuable as a warm start [61], as it encodes transferable geometric inductive biases that can accelerate optimization or reduce data requirements[62].
@@ Conclusion
In this paper, we introduced HiLAB, a novel hybrid inverse-design framework that effectively combines partial TO, generative AI via a ViT-based VAE, and BO to efficiently address complex multifunctional photonic device design challenges. By truncating traditional adjoint-driven TO at early stages and systematically varying key physical parameters, we generate diverse and computationally affordable datasets that capture the critical structural features necessary for high-performance device design. The trained VAE encodes these partially optimized structures into a compact latent space, significantly reducing computational demands for subsequent design exploration. By leveraging BO within this learned latent space, HiLAB systematically guides the search towards global optima (or strong local optima) while simultaneously optimizing physical hyperparameters such as device thickness and period. We demonstrated the effectiveness of HiLAB by designing an achromatic beam deflector achieving balanced diffraction efficiencies, theoretically exceeding 24% across red, green, and blue wavelengths—substantially outperforming similar devices designed by prior conventional TO methods and significantly reducing chromatic aberrations. Crucially, the generality and flexibility of HiLAB mean it can be easily adapted for different design targets, constraints, or fabrication tolerances without necessitating extensive additional simulations or retraining. By providing a systematic, scalable, and efficient inverse-design methodology, HiLAB holds considerable promise for accelerating the development of robust, high-performance photonic and metamaterial devices across diverse application domains.